# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ Appendix A Appendix
We begin by proving Proposition 3.1 which is needed to demonstrate the existence of the fusion systems introduced in Section 3.
Proposition A.1 (Proposition 3.1).
The map $\psi^{*}:N_{P^{*}}(R)\rightarrow P_{R}^{*}$ given by
is a monomorphism. Furthermore, the following hold:
  1. (1)
    $R\psi^{*}=O_{p}(P_{R})$; and
  2. (2)
Proof.
$=\left(\phi\Phi,\theta^{p^{l}}\Pi,\left(\begin{matrix}a^{p^{l}}&0\\ c^{p^{l}}&b^{p^{l}}\end{matrix}\right)\left(\begin{matrix}\alpha&0\\ \kappa&\beta\end{matrix}\right),\left(\lambda^{p^{l}}\overline{y^{p}}+\mu^{p^{l}}\overline{xy^{p-1}}+\nu^{p^{l}}\overline{x^{2}y^{p-2}}\right)\cdot\left(\Pi,\left(\begin{matrix}\alpha&0\\ \kappa&\beta\end{matrix}\right)\right)+\chi\overline{y^{p}}+\tau\overline{xy^{p-1}}+\eta\overline{x^{2}y^{p-2}}\right)$
$=\left(\phi\Phi,\theta^{p^{l}}\Pi,\left(\begin{matrix}a^{p^{l}}\alpha&0\\ c^{p^{l}}\alpha+\kappa b^{p^{l}}&b^{p^{l}}\beta\end{matrix}\right),(\lambda^{p^{l}}\Pi^{-1}\beta^{-p}+\chi)\overline{y^{p}}+(\nu^{p^{l}}\Pi^{-1}\kappa\alpha^{-2}\beta^{1-p}+\mu^{p^{l}}\Pi^{-1}\alpha^{-1}\beta^{1-p}+\tau)\overline{xy^{p-1}}+(\nu^{p^{l}}\Pi^{-1}\alpha^{-2}\beta^{2-p}+\eta)\overline{x^{2}y^{p-2}}\right).$
On the other hand,
$=\left(\phi\Phi,\left(\begin{smallmatrix}\theta(ab^{p})^{p^{l}}&0&0&(\theta ab^{p}\lambda)^{p^{l}}\\ 0&(\theta a^{2}b^{p-1})^{p^{l}}&(\theta a^{2}b^{p-1}\nu)^{p^{l}}&(\theta a^{2}b^{p-1}\mu)^{p^{l}}\\ 0&0&b^{p^{l}}&c^{p^{l}}\\ 0&0&0&a^{p^{l}}\end{smallmatrix}\right)\left(\begin{smallmatrix}\Pi\alpha\beta^{p}&0&0&\Pi\alpha\beta^{p}\chi\\ 0&\Pi\alpha^{2}\beta^{p-1}&\Pi\alpha^{2}\beta^{p-1}\eta&\Pi\alpha^{2}\beta^{p-1}\tau\\ 0&0&\beta&\kappa\\ 0&0&0&\alpha\end{smallmatrix}\right)\right)$
$=\left(\phi\Phi,\left(\begin{smallmatrix}(\theta ab^{p})^{p^{l}}\Pi\alpha\beta^{p}&0&0&(\theta ab^{p})^{p^{l}}\Pi\alpha\beta^{p}\chi+(\theta ab^{p}\lambda)^{p^{l}}\alpha\\ 0&(\theta a^{2}b^{p-1})^{p^{l}}\Pi\alpha^{2}\beta^{p-1})&(\theta a^{2}b^{p-1})^{p^{l}}\Pi\alpha^{2}\beta^{p-1}\eta+(\theta a^{2}b^{p-1}\nu)^{p^{l}}\beta)&(\theta a^{2}b^{p-1})^{p^{l}}\Pi\alpha^{2}\beta^{p-1}\tau+(\theta a^{2}b^{p-1}\nu)^{p^{l}}\kappa+(\theta a^{2}b^{p-1}\mu)^{p^{l}}\alpha)\\ 0&0&b^{p^{l}}\beta&b^{p^{l}}\kappa+\alpha c^{p^{l}}\\ 0&0&0&a^{p^{l}}\alpha\end{smallmatrix}\right)\right).$
It is elementary to check that (1) and (2) hold. ∎
Next we prove some group and module recognition results which, while not needed in the proofs of our main results, may still be of interest.
@@
Let $S$ and $V$ be as described in the proposition. Note that each $s\in S\setminus V$ acts cubically on $V$. Indeed, we have that $[V,s]W=[V,S]W$ and $[[V,S]W,s]=C_{W}(S)$ for all such $s$. It follows that $[V,S]W=Z_{2}(S)$ has order $q^{3}$.
Assume that $E\in\mathcal{E}(\mathcal{F})\setminus\{V\}$ so that $Z_{2}(S)\leq N_{S}(E)$. Suppose that $\Omega(Z(E))\not\leq V$. Then $E\cap V\leq C_{V}(\Omega(Z(E)))$ has order at most $q^{2}$ so that $|Z_{2}(S)E/E|\geqslant q$. An application of Lemma 5.5 reveals that $S=\Omega(Z(E))V$, $E=\Omega(Z(E))$ and $Z_{2}(S)E=N_{S}(E)$. Moreover, $E\cap V=Z(S)$ has order $q^{2}$. If $S$ was quadratic on $V$ then since $S=EV$, we must have that $[E,V]\leq C_{V}(E)=E\cap V=Z(S)$ from which we conclude that $E\trianglelefteq S$, a contradiction. Hence, the proposition holds in this case.
Suppose now that $\Omega(Z(E))\leq V$. Then $[Z_{2}(S),E]\leq Z(S)\leq\Omega(Z(E))$ and an application of Lemma 6.4 yields that $Z_{2}(S)\leq E$. Then $E\trianglelefteq S$ and $Z_{2}(S)$ is an elementary abelian subgroup of $E$ of order $q^{3}$. Let $A$ be an elementary abelian subgroup of $E$ of order at least $q^{3}$. Assume that $A\not\leq V$. Then $q^{3}\leqslant|A|\leqslant|AV/V||C_{V}(A)|\leqslant q.q^{2}$ and so $|A|=q^{3}$. Suppose that every elementary abelian subgroup of $E$ of maximal order has order $q^{3}$, and there exists such an $A$ lying outside of $V$. In particular, $E\cap V=Z_{2}(S)$. Then $A\cap V$ is centralised by $S=AV$ and so $Z(S)=Z(E)$ has order $q^{2}$. If $S$ acted quadratically on $V$, $S^{\prime}=[A,V]\leq C_{A}(V)=Z(S)$ and Lemma 6.4 implies that $S\leq E$, a contradiction.
@@
Write ¯$\overline{\mathcal{F}}:=\mathcal{F}/Z$. By [AKO11, Theorem I.3.10] it suffices to show that $\mathcal{F}$ is $\mathcal{F}^{cr}$-saturated. If $P\in\mathcal{F}^{cr}$ then $\overline{P}\in\overline{\mathcal{F}}^{cr}$ by [Cra11, Proposition 5.60] and since $\overline{\mathcal{F}}$ is ¯$\overline{\mathcal{F}}^{cr}$-generated, $\overline{P}$ is conjugate to a fully $\overline{\mathcal{F}}$-automised, $\overline{\mathcal{F}}$-receptive subgroup $\overline{Y}$. Then $\overline{Y}$ is fully $\mathcal{F}$-normalised by [AKO11, Lemma I.2.6] so that $|N_{\overline{S}}(\overline{Y})|\geqslant|N_{\overline{S}}(\overline{A})|$ for all $\overline{A}\in\overline{Y}^{\overline{\mathcal{F}}}$. Then for any $B\leq S$ with $Z\leq B$, we have that $|N_{S}(B)/Z|=|N_{\overline{S}}(\overline{B})|$ and we deduce that $|N_{S}(Y)/Z|\geqslant|N_{S}(A)/Z|$. Therefore, $|N_{S}(Y)|\geqslant|N_{S}(A)|$ and $Y$ is fully $\mathcal{F}$-normalised. Applying [AKO11, Lemma I.2.5] yields that $Y$ is fully $\mathcal{F}$-automised.

### v2

@@ Appendix A Appendix
We begin this appendix with some group and module recognition results which, while not needed in the proofs of our main results, may still be of interest.
@@
Let $S$ and $V$ be as described in the proposition. Note that each $s\in S\setminus V$ acts cubically on $V$. Indeed, we have that $[V,s]W=[V,S]W$ and $V,S]W,s]=C_{W}(S)$ for all such $s$. It follows that $V,S]W=Z_{2}(S)$ has order $q^{3}$.
Assume that $E\in\mathcal{E}(\mathcal{F})\setminus\{V\}$ so that $Z_{2}(S)\leq N_{S}(E)$. Suppose that $\Omega(Z(E))\not\leq V$. Then $E\cap V\leq C_{V}(\Omega(Z(E)))$ has order at most $q^{2}$ so that $|Z_{2}(S)E/E|\geqslant q$. An application of Lemma 5.5 reveals that $S=\Omega(Z(E))V$, $E=\Omega(Z(E))$ and $Z_{2}(S)E=N_{S}(E)$. Moreover, $E\cap V=Z(S)$ has order $q^{2}$. If $S$ was quadratic on $V$ then since $S=EV$, we must have that $E,V]\leq C_{V}(E)=E\cap V=Z(S)$ from which we conclude that $E\trianglelefteq S$, a contradiction. Hence, the proposition holds in this case.
Suppose now that $\Omega(Z(E))\leq V$. Then $Z_{2}(S),E]\leq Z(S)\leq\Omega(Z(E))$ and an application of Lemma 6.4 yields that $Z_{2}(S)\leq E$. Then $E\trianglelefteq S$ and $Z_{2}(S)$ is an elementary abelian subgroup of $E$ of order $q^{3}$. Let $A$ be an elementary abelian subgroup of $E$ of order at least $q^{3}$. Assume that $A\not\leq V$. Then $q^{3}\leqslant|A|\leqslant|AV/V||C_{V}(A)|\leqslant q.q^{2}$ and so $|A|=q^{3}$. Suppose that every elementary abelian subgroup of $E$ of maximal order has order $q^{3}$, and there exists such an $A$ lying outside of $V$. In particular, $E\cap V=Z_{2}(S)$. Then $A\cap V$ is centralised by $S=AV$ and so $Z(S)=Z(E)$ has order $q^{2}$. If $S$ acted quadratically on $V$, $S^{\prime}=[A,V]\leq C_{A}(V)=Z(S)$ and Lemma 6.4 implies that $S\leq E$, a contradiction.
@@
Write $\overline{\mathcal{F}}:=\mathcal{F}/Z$. By [AKO11, Theorem I.3.10] it suffices to show that $\mathcal{F}$ is $\mathcal{F}^{cr}$-saturated. If $P\in\mathcal{F}^{cr}$ then $\overline{P}\in\overline{\mathcal{F}}^{cr}$ by [CRA11, Proposition 5.60] and since $\overline{\mathcal{F}}$ is $\overline{\mathcal{F}}^{cr}$-generated, $\overline{P}$ is conjugate to a fully $\overline{\mathcal{F}}$-automised, $\overline{\mathcal{F}}$-receptive subgroup $\overline{Y}$. Then $\overline{Y}$ is fully $\mathcal{F}$-normalised by [AKO11, Lemma I.2.6] so that $|N_{\overline{S}}(\overline{Y})|\geqslant|N_{\overline{S}}(\overline{A})|$ for all $\overline{A}\in\overline{Y}^{\overline{\mathcal{F}}}$. Then for any $B\leq S$ with $Z\leq B$, we have that $|N_{S}(B)/Z|=|N_{\overline{S}}(\overline{B})|$ and we deduce that $|N_{S}(Y)/Z|\geqslant|N_{S}(A)/Z|$. Therefore, $|N_{S}(Y)|\geqslant|N_{S}(A)|$ and $Y$ is fully $\mathcal{F}$-normalised. Applying [AKO11, Lemma I.2.5] yields that $Y$ is fully $\mathcal{F}$-automised.